OpenStudy (anonymous):

Help please. Will give medal. I need to convert 12.71 (g) of NcHCO3 to NaCl ?

OpenStudy (aaronq):

you can't convert elements to other elements

OpenStudy (anonymous):

I don't know but it is in my chemistry homework and it says and the teacher said to convert the above 12.71 g of NcHCO3 to NaCI

OpenStudy (aaronq):

was an equation provided for the process? NaCl is a salt, while NaHCO3 can act as a base or an acid. what are the other reagents?

OpenStudy (anonymous):

no equation provided

OpenStudy (anonymous):

This info is all i was given

OpenStudy (aaronq):

is it part of another question? it doesn't make sense

OpenStudy (anonymous):

This is the actual question. The NaHCO3 is the limiting reactant and the HCl is the excess reactant in this experiment. Determine the theoretical yield of the NaCl product, showing all of your work in the space below. This was my answer: NaHCO3 – The empty dish 37.06 (g) - 24.35 (g) =12.71 (g) of NcHCO3 ** Then the teacher told me i needed to convert 12.71 g of NcHCO3 to NaCI

OpenStudy (aaronq):

the other reactant is HCl which is in excess. thus the limiting reactant is \(NaHCO_3\). so you have \(NaHCO_3 +HCl\rightarrow H_2O+NaCl\) it needs to be balanced. Are you comfortable with stoichiometry problems?

OpenStudy (aaronq):

\(NaHCO_3+HCl→H_2O+NaCl + CO_2\) i forgot it makes CO2 as well

OpenStudy (anonymous):

not so much chemistry is not my best subject

OpenStudy (aaronq):

okay, once you have the balanced equation, convert the mass of NaHCO3 to moles (use equation at the bottom). Then use the stoichiometric coefficients to find moles produced. Set up a ratio using the species of interest, like so: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients , \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) From here you can isolate what you need. For example: if you have 2 moles of B, how many moles of C can you produce? solve algebraically: \(\dfrac{2}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\rightarrow n_C=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\) -------------------------------------------------------- To convert mass to moles, use the relationship: \(n=\dfrac{m}{M}\) where, M=molar mass, m=mass, and n= moles. --------------------------------------------------

OpenStudy (anonymous):

Thank you so much i will see if that helps

OpenStudy (aaronq):

no problem, good luck